COHERENT EFFECTS INVOLVING THE EXCITATION AND RELAXATION OF THE COUPLED L_{a}/L_{b} ELECTRONIC EXCITED STATES OF NAPHTHALENE: A TIME DEPENDENT EXPERIMENTAL STUDY

\underline{R. MONTERO}, A. PERALTA CONDE, F. CASTANO, AND A. LONGARTE, Departamento de Química-Física, Facultad de Ciencia y Tecnología, Universidad del País Vasco, Apartado 644, ES-48080 Bilbao, Spain.

The ultrafast dynamics of the non-adiabatic coupling between the L_{a}(S_{2})/L_{b}(S_{1}) electronic excited states of naphthalene has been investigated, regarding the coherent aspects of the preparation and subsequent relaxation of the system. This electronic coupling represents a well known case of non-adiabatic behavior that has been used for years as a benchmark to test theoretical models. The jet cooled naphthalene molecule was prepared in the L_{a} and L_{b} states by femtosecond pump pulses in the UV region (318-268 nm), while the temporal evolution of the system was tracked by multiphoton ionization of the molecule with probe pulses centered at 800 nm. The time dependent signals collected at excitation wavelengths corresponding to the weak S_{0}-L_{b} transition are dominated by the CPR (Coherent Population Transfer) effect induced by the blue-shifted stronger S_{0}-L_{a} absorption. The CPR effect results in the transient population of the L_{a} state during the interaction of the pump pulse with the system, revealing the coherent nature of the excitation process. The transients collected after excitation to the L_{a} state reveal the ultrafast relaxation to the strongly coupled L_{b} state. The electronic and/or vibrational nature of the periodic recurrences observed along the relaxation of the electronic population will be discussed.